OpenStudy (anonymous):

Find all the zeros of each equation. 7x^2 – 144 = –x^4 3, –3, 4i, 0 3, 4i 3, –3, 4i, –4i –3, –4i

OpenStudy (anonymous):

what ?

hero (hero):

x^4 + 7x^2 - 144 = 0 x^4 + 16x^2 - 9x^2 - 144 = 0 x^2(x^2 + 16) - 9(x^2 + 16) = 0 (x^2 + 16) (x^2 - 9) = 0 x^2 + 16 = 0 x^2 - 9 = 0 x^2 = -16 x^2 = 9 x = + or - 4i x = + or - 3 So the four solutions are: x = {-4i, 4i, -3, 3}

OpenStudy (anonymous):

thank you soo much , i appreciate it ! (: